# SurGE: A Benchmark and Evaluation Framework for Scientific Survey Generation

### v2

@@ D.3 StepSurvey
StepSurvey (Lai et al., 2024b) is a more granular generation strategy that also begins with retrieving the top 100 candidate papers for a given topic $t$ but proceeds through distinct planning and drafting phases in a sequential manner. This baseline is proposed by a team named “ID” in the NLPCC2024 competition task 6 (Tian et al., 2024b). Rather than producing an overarching outline at once, it starts by proposing a survey title and a set of primary headings that collectively capture the central themes of the retrieved literature. Subsequently, it uses the primary headings to guide the selection of secondary or finer-grained topics, each mapped to a relevant subset of the retrieved papers. The LLM then produces a full draft by writing each subsection with explicit attention to references and academic conventions, thereby encouraging greater control and consistency across sections. Throughout this process, the system attempts to maintain a balanced level of detail, striving for a clear exposition of important subtopics while avoiding excessive verbosity or redundancy. By structuring the content in incremental steps, StepSurvey aims to achieve coherent organization and thorough coverage of the literature.
Appendix E Prompt Template
@@
Appendix F Selected Ground Truth Survey
@@
Automated Machine Learning on Graphs: A Survey 2021 cs.LG 97
[…]
A Survey on Multi-modal Summarization 2021 cs.CL 73
Towards Reasoning in Large Language Models: A Survey 2022 cs.CL 841
Creativity and Machine Learning: A Survey 2021 cs.LG 56
@@
Transformers in Medical Imaging: A Survey 2022 eess.IV 889
A Survey of Privacy Attacks in Machine Learning 2020 cs.CR 368
Survey on Evaluation Methods for Dialogue Systems 2019 cs.CL 394

### v4

@@ E.3 StepSurvey
StepSurvey (Lai et al., 2024) is a more granular generation strategy that also begins with retrieving the top 100 candidate papers for a given topic $t$ but proceeds through distinct planning and drafting phases sequentially. This baseline is proposed by a team named “ID” in the NLPCC2024 competition task 6 (Tian et al., 2024b). Rather than producing an overarching outline at once, it starts by proposing a survey title and a set of primary headings that collectively capture the central themes of the retrieved literature. Subsequently, it uses the primary headings to guide the selection of secondary or finer-grained topics, each mapped to a relevant subset of the retrieved papers. The LLM then produces a full draft by writing each subsection with explicit attention to references and academic conventions, thereby encouraging consistency across sections. Throughout this process, the system attempts to maintain a balanced level of detail, striving for a clear exposition of important subtopics while avoiding excessive verbosity or redundancy. By structuring the content in incremental steps, StepSurvey aims to achieve coherent organization and thorough coverage of the literature.
E.4 SurveyForge
SurveyForge (Yan et al., 2025) introduces a two-stage framework that combines heuristic outline generation with memory-driven content writing. To ensure a fair comparison, we strictly follow the official implementation [^3] https://github.com/InternScience/SurveyForge but replace the original retrieval component with the retriever described in Section 5.1, keeping the candidate document pool the same as other baselines. The pipeline begins by constructing a hierarchical outline, leveraging structural patterns derived from human-authored surveys to organize the topics. Subsequently, the method employs a Scholar Navigation Agent (SANA) to synthesize the content. In our adapted workflow, this agent queries the pre-retrieved papers to draft specific subsections, ensuring that the generated text is structurally coherent and grounded in the provided references.
Appendix F Extended Related Work
F.1 Retrieval-Augmented Generation
Retrieval-Augmented Generation (RAG) has emerged as a cornerstone for scientific writing (Lewis et al., 2020; Gao et al., 2023; Su et al., 2024). Fundamentally, RAG pipelines follow a retrieve-then-read paradigm, identifying relevant documents to condition the generator (Guu et al., 2020). While this paradigm shares the same core logic as Survey Generation (incorporating external knowledge to generate content), standard RAG approaches struggle to meet the unique demands of survey writing. Specifically, naive RAG pipelines face three critical limitations in this context. First, standard RAG pipelines lack the hierarchical planning required to maintain a logical narrative flow, often treating the task as simple QA rather than document composition. Second, they fail to capture complex inter-document relationships (e.g., chronological evolution or methodological contrasts), which are essential for synthesizing conflicting information from massive corpora. Third, the limited context window in standard models often leads to information loss when processing hundreds of references. To address the structural complexity of full-length surveys, the field is shifting towards specialized Agentic Workflows that decompose the task into autonomous phases of planning, execution, and verification. We introduce this emerging paradigm in the following subsection.
F.2 Agentic Survey Generation
Before the advent of Large Language Models, scientific survey generation was often viewed as an advanced form of Multi-Document Summarization (MDS), yet it presents unique challenges regarding scale and structure Lu et al. (2020); Ma et al. (2022). Unlike traditional MDS, which focuses on distilling information into concise overviews (Nayeem and Rafiei, 2024; Bleiweiss, 2023), survey generation aims to synthesize extensive literature into a coherent, lengthy manuscript with a hierarchical structure Wang et al. (2024).
Early approaches struggled with these constraints due to the limited context windows of language models. The advent of LLMs has triggered a paradigm shift from simple multi-document summarization to complex agentic workflows. For instance, AutoSurvey (Wang et al., 2024) pioneers a multi-stage pipeline that moves beyond simple drafting. It generates a structured outline and performs iterative expansions, explicitly refining sections to address factual and stylistic inconsistencies. Similarly, StepSurvey (Lai et al., 2024) employs a granular strategy, prioritizing sequential planning where secondary topics are mapped to specific paper subsets before drafting to ensure balanced detail. From another angle, SurveyForge (Yan et al., 2025) introduces a distinct approach by combining heuristic outline generation with a memory-driven Scholar Navigation Agent (SANA) to synthesize content. Most recently, SurveyX (Liang et al., 2025) introduces AttributeTree templates to distill reference information, and uses extraction and generation templates to automatically construct auxiliary figures and tables, followed by a re-polishing (post-refinement) process.
Despite these architectural innovations, evaluation remains a critical bottleneck. Existing works predominantly operate within a "self-validation" paradigm, where metrics are either unscalable or tailored to specific pipelines. For instance, approaches like StepSurvey rely heavily on human evaluation (Tian et al., 2024a), which strictly limits scalability and reproducibility. Conversely, framework-oriented works such as AutoSurvey and SurveyForge propose custom metrics designed to validate their own specific workflows. This fragmentation creates a "metric bias" that obscures fair cross-system comparison, highlighting the urgent need for a standardized, method-agnostic benchmark to assess these diverse agentic strategies.
F.3 Long-Context Evaluation and Long-Form Generation
[…]
In parallel, long-form generation has been widely studied under long-document summarization and long-form QA. Datasets such as GovReport Huang et al. (2021) and QMSum Zhong et al. (2021) require generating summaries from lengthy, information-dense sources, and BookSum Kryściński et al. (2022) focuses on narrative summarization with rich discourse and long-range dependencies. Despite their value, these tasks typically target a summary or answer and rarely require explicit bibliographic grounding, faithful claim-to-citation mapping, or field-level structural synthesis.
Thus, scientific survey generation raises a different evaluation challenge. A survey is not only long, but it must also be academically valid, requiring (i) broad bibliographic coverage of seminal work, (ii) faithful attribution of claims to references, and (iii) coherent structural taxonomy that organizes and contrasts prior studies. This gap motivates a dedicated, method-agnostic benchmark and diagnostic evaluation framework that can fairly compare diverse survey generation pipelines.
Appendix G Prompt Template
@@
Appendix H Selected Ground Truth Survey
@@
Automated Machine Learning on Graphs: A Survey 2021 cs.LG 97
[…]
A Survey on Multi-modal Summarization 2021 cs.CL 73
Towards Reasoning in Large Language Models: A Survey 2022 cs.CL 841
Creativity and Machine Learning: A Survey 2021 cs.LG 56
@@
Transformers in Medical Imaging: A Survey 2022 eess.IV 889
A Survey of Privacy Attacks in Machine Learning 2020 cs.CR 368
Survey on Evaluation Methods for Dialogue Systems 2019 cs.CL 394